Acme Company’s production budget for August is 19,000 units and includes the following component unit costs: direct materials, $8.0; direct labor, $11.6; variable overhead, $5.6. Budgeted fixed overhead is $47,000. Actual production in August was 20,445 units. Actual unit component costs incurred during August include direct materials, $9.80; direct labor, $10.00; variable overhead, $6.40. Actual fixed overhead was $50,000. The standard variable overhead rate per unit consists of $5.6 per machine hour and each unit is allowed a standard of 1 hour of machine time. During August, $130,848 of actual variable overhead cost was incurred for 22,560 machine hours.

Required:
Calculate the variable overhead spending variance and the variable overhead efficiency variance.
Answer #1

Solution:

Standard rate of variable overhead = $5.60 per machine hour

Actual rate of variable overhead = $130,848 / 22560 = $5.80 per machine hour

Standard hours for actual production = 20445*1 = 20445 hours

Variable overhead spending variance = (SR - AR) * AH = ($5.60 - $5.80) * 22560 = $4,512 U

Variable overhead efficiency variance = (SH - AH) * SR = (20445 - 22560) * $5.60 = $11,844 U